OpenStudy (anonymous):

Find the inverse of the following function: f(x)=2^(x-1)+3

OpenStudy (anonymous):

ln(y/6)+1=x

OpenStudy (anonymous):

ln both sides, then solve for x

OpenStudy (nikvist):

\[f(x)=2^{x-1}+3\]\[f(x)-3=2^{x-1}\]\[\log_2{(f(x)-3)}=x-1\]\[1+\log_2{(f(x)-3)}=x\]\[\log_2{(2f(x)-6)}=x\]\[\Rightarrow\quad f^{-1}(x)=\log_2{(2x-6)}\]